OpenStudy (anonymous):

solve sin(3x)=1 on the interval [0,(2pi\3)]

OpenStudy (anonymous):

this means \[3x=0\] or \[3x=\pi\] or \[3x=2\pi\] etc

OpenStudy (anonymous):

because those are the inputs that give 0 for sine. so divide by 3 to get \[x=0\] or \[x=\frac{\pi}{3}\]or \[x=\frac{2\pi}{3}\]

OpenStudy (anonymous):

it doesnt give me that option

OpenStudy (anonymous):

that is because i am a moron. it says = 1 not = 0 lets try again

OpenStudy (anonymous):

this means \[3x=\frac{\pi}{2}\] pr \[3x=\frac{3\pi}{2}\] or \[3x=\frac{5\pi}{2}\] etc

OpenStudy (anonymous):

and so \[x=\frac{\pi}{6}\] or \[x=\frac{\pi}{3}\] or \[x=\frac{5\pi}{6}\] etc

OpenStudy (anonymous):

man i must be tired. second one should be \[x=\frac{\pi}{2}\]